A sugar solution was made by mixing 8 qt of a 2% sugar solution and 6 qt of a 51% sugar solution. find the concentration of the new mixture
Answer by stanbon(75887)

A sugar solution was made by mixing 8 qt of a 2% sugar solution and 6 qt of a 51% sugar solution. find the concentration of the new mixture
-------
Equation:
sugar + sugar = sugar
0.02*8 + 0.51*6 = x(8+6)
-----
Multiply thru by 100 to get:
-----
2*8 + 51*6 = 1400x
-----
16 + 306 = 1400x
322 = 1400x
x = 322/1400
x = 0.23 = 23% (That is the concentration of the mixture)
==================
Cheers,
Stan H.
==================